Suppose the mean number of days to germination of a variety of seeds is 22, with a standard deviation of 2.3 days. Find the probability that the mean germination time of a sample of 160 seeds will be within 0.5 days of the population mean.


1
Expert's answer
2022-06-01T14:20:18-0400
"P(-\dfrac{0.5}{2.3/\sqrt{160}}<Z<\dfrac{0.5}{2.3/\sqrt{160}})"

"\approx P(Z<2.7498)-P(Z\le-2.7498)"

"\approx0.99702-0.00298=0.9940"